OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/4x + 2. y = 4x - 1 y = 4x - 3 y = x - 3 y = x - 1

OpenStudy (anonymous):

Are those supposed to be my options?

OpenStudy (anonymous):

yep !

OpenStudy (anonymous):

Your problem is broken.

OpenStudy (anonymous):

wait no .. i typed the last 2 wrong .. y = 1/4x - 3 y = 1/4x - 1

OpenStudy (anonymous):

Okay, let me ask you about what you already know. Do you know slope intercept form of an equation? Do you know what makes two lines parallel?

OpenStudy (anonymous):

y=mx+b ? and when 2 lines remain constant with their distance .. right ?